# C-DiffSET: Leveraging Latent Diffusion for SAR-to-EO Image Translation with Confidence-Guided Reliable Object Generation

### v2

@@ Abstract
Figure 1: Qualitative results of SAR-to-EO image translation using our proposed C-DiffSET framework (Confidence Diffusion for SAR-to-EO Translation). Rows 1, 2, and 3 display translations on the QXS-SAROPT, SAR2Opt, and SpaceNet6 datasets, respectively. Our proposed C-DiffSET achieve superior structural accuracy and visual fidelity compared to very recent methods.
@@ 1 Introduction
Satellite imagery plays a critical role in various applications, including surveillance, transportation, agriculture, disaster assessment, and environmental monitoring [81, 60, 70, 4, 67]. A significant portion of these applications relies on Electro-Optical (EO) imagery, which provides multi-spectral data captured by EO sensors. However, a fundamental limitation of EO imagery is its susceptibility to weather and lighting conditions, reducing its usability in scenarios such as cloudy weather or nighttime. In contrast, Synthetic Aperture Radar (SAR) offers robust sensing capabilities under all weather conditions without the need for light, making it ideal for nighttime operations. Despite these advantages, SAR imagery presents several challenges. SAR data suffers from speckle noise due to the backscattering of electromagnetic waves [13, 8, 79, 57]. Additionally, SAR images are typically grayscale, lacking the rich spectral and color information in EO images, making them difficult to be interpreted, especially for non-experts. To enhance the usability of SAR data, SAR-to-EO image translation (SET) has emerged as a promising research area in remote sensing, facilitating scene interpretation for tasks such as cloud removal [52, 41, 12] and change detection [39, 40, 38].
@@
Recent advancements [75, 50, 25, 31, 32, 54, 14, 28] have demonstrated the potential of generating EO-like outputs from SAR images, offering improved textures and enhanced color information. However, several challenges persist: (i) The domain gap between SAR and EO images makes SET an ill-posed problem, exacerbated by the inherent noise in SAR images; (ii) The misalignment between SAR and EO datasets is prone to occur due to their different sensor platforms, satellite positioning shifts, or acquisition conditions, as shown in Fig. 2; (iii) Temporal misalignments (discrepancies) occur due to different acquisition times, also resulting in the spatial misalignment with differences in object appearance and seasons. This makes it very complicated the SET problem (Fig. 2); (iv) Finally, the scarcity of paired SAR-EO datasets makes it challenging for existing methods to learn the SET effectively, often leading to overfitting or unstable results.
@@
To address these limitations, we firstly propose a novel framework that leverages a pretrained Latent Diffusion Model (LDM) [47, 11] to improve SAR-to-EO image translation, which is denoted as Confidence Diffusion for SAR-to-EO Translation (C-DiffSET). The advantages of utilizing the pretrained LDM for SET tasks are as follows: (i) Since EO images share visual characteristics with natural images, we fine-tune the pretrained LDM—trained on large-scale natural image datasets—to transfer its representation power to the SET task. This addresses the issue of limited SAR-EO paired data, as our framework can leverage the pretrained knowledge from natural image distributions; (ii) The pretrained LDM operates in a down-sampled latent space rather than the pixel domain, making it more robust to local spatial misalignments caused by imperfect alignment processes; (iii) The variational auto-encoder (VAE) [29] from LDM can also be applicable for SAR images. As shown in Fig. 4, despite the presence of significant speckle noise in SAR data, the VAE encoder effectively embeds SAR images into the same latent space as EO images, leveraging its denoising capability. This enables the smooth transfer of SAR latents as conditioning information in the reverse diffusion process, ensuring the generation of EO outputs with accurate pixel-wise correspondence to the SAR inputs.
@@
Figure 2: Examples of misalignments in paired SAR-EO datasets. Left: Local spatial misalignments caused by sensor differences or acquisition conditions. Right: Temporal misalignments (discrepancies) where objects (e.g., ships) appear or disappear between SAR and EO images due to their different acquisition times.
@@
In addition to utilizing the pretrained LDM as a foundation model for our SET task, we introduce a confidence-guided diffusion (C-Diff) loss to handle temporal discrepancies. The temporal discrepancies, such as objects appearing in only one modality, can introduce artifacts and hallucinated content in the generated EO images. To mitigate this, the U-Net [48] in our framework predicts both each predicted noise and its corresponding confidence map, which quantifies the pixel-wise uncertainty in the predicted noise. The U-Net receives a channel-wise concatenated SAR and noisy EO features as input, allowing it to jointly model both contexts. This setup enables the confidence map to reflect the inherent uncertainty caused by temporal discrepancies, guiding the C-Diff loss to adaptively reduce penalties in regions where SAR-EO misalignments likely occur. As a result, the generated EO images achieve high pixel-wise fidelity while minimizing artifacts and hallucinations. Our key contributions are summarized as:
@@
  • We propose C-DiffSET, the first framework to fine-tune a pretrained LDM for SET tasks, effectively leveraging their learned representations to overcome the scarcity of SAR-EO pairs and local misalignment challenges.
@@
  • We introduce a novel C-Diff loss that can guide our C-DiffSET to reliably predict both EO outputs and confidence maps for accurate SET.
@@ 2.1 Image-to-Image Translation
Image-to-image translation has been widely studied in various fields, such as image colorization [76, 77], style transfer [84, 24, 55], and image inpainting [22, 43]. These models typically rely on carefully constructed paired datasets across domains for effective training. For instance, Pix2Pix [24] introduced conditional GANs with an $l_{1}$ loss for domain-specific translations, while ASAPNet [55] proposed a lightweight framework for converting segmentation maps into realistic images, demonstrating abstract-to-detailed generation. However, acquiring paired datasets remains challenging, and various methods have been proposed to work around unpaired data constraints [1, 35, 27]. CycleGAN [84] addressed this limitation using a cycle-consistency loss to facilitate training without paired data. StarGAN [7] extended this idea to support multiple domain mappings within a single model, though its deterministic one-to-one approach limits flexibility. MUNIT [20] overcame this by decomposing images into domain-invariant content and domain-specific style representations, enabling diverse mappings while maintaining content integrity. Further addressing non-bijective translation, StegoGAN [69] introduced steganography into GAN-based models, enhancing semantic consistency in cases of domain mismatch by reducing spurious features in generated images without additional supervision.
Diffusion-based approaches. Diffusion models [18, 47, 11, 44, 17, 49, 33], have gained significant attraction in image-to-image translation due to their iterative denoising process, effectively modeling complex data distributions. Palette [49] utilizes conditional Denoising Diffusion Probabilistic Models (DDPMs) [18] in the pixel domain, achieving high-quality translations. Pixel-space operations often incur high computational costs for high-resolution images, motivating recent advances that utilize latent space diffusion to achieve greater efficiency. BBDM [33] diffuses directly from input to target latent space, eliminating the need for separate noise conditioning but imposing strict pixel-wise alignment requirements between input and target images. DGDM [71] is built on BBDM with a deterministic approach, where a translator network generates target features prior to diffusion, adding computational overhead and dependency on the initial translator’s performance.
@@ 2.2 SAR-to-EO Image Translation (SET)
Deep learning-based image-to-image translation methods [24, 55, 22, 20, 1, 35, 82, 33] have been widely adopted to remote sensing applications, including SAR-to-EO image translation (SET). However, SAR imagery poses unique challenges due to inherent characteristics like speckle noise from backscattering effects [13, 8, 79, 57, 59, 74, 21], complicating the translation process. To address these challenges, most SET studies employ GAN-based approaches [63, 5, 66, 32, 30, 9, 75, 72]. Conditional GAN-based models are commonly used to leverage noisy SAR information as a conditioning input, while CycleGAN-based models align the SAR and EO domains by enforcing cycle-consistency constraints. MCGAN [25] utilized a CycleGAN-based model, incorporating a mask vector with class labels to represent different regions. DTCDN [34] employed NICE-GAN [6] for joint SET and change detection tasks, improving feature alignment between domains. PLFM [52] introduced a multimodal model that integrates spatial and temporal data with SAR guidance images to produce clear EO outputs. Additionally, CFCA-SET [31] proposed a coarse-to-fine SAR-to-EO model that incorporates near-infrared (NIR) images during training to refine EO generation.
@@
Despite these advancements, both GAN-based and diffusion-based methods encounter limitations stemming from the scarcity of paired SAR-EO datasets. GAN-based models frequently face with convergence challenges and mode collapse under limited data, while diffusion-based approaches struggle with overfitting and slow convergence due to the need for extensive training from scratch. To overcome these constraints, our C-DiffSET leverages a pretrained LDM as foundational model. By capitalizing on the extensive generative priors of large-scale text-to-image models, the C-DiffSET effectively adapts to the SET task, achieving robust translation with a reduced overfitting risk.
2.3 Latent Diffusion Models (LDMs)
Latent Diffusion Models (LDMs) [47, 11] extend the foundational DDPMs [18], which apply iterative noise addition to pixel-space images in a forward process, transforming them toward following Gaussian distribution. DDPMs are trained to reverse this process, step-by-step denoising to recover high-quality images. While effective for complex data modeling, DDPMs are computationally demanding in high-resolution applications due to their pixel-space processing. In contrast, LDMs address this by shifting the diffusion process to a learned latent space, significantly reducing memory requirements and computational costs. Using a VAE to encode images into a compact latent space, LDMs perform denoising directly on these representations, retaining high-resolution details with improved efficiency.
LDMs as foundation models. Open-source [62] releases of foundation LDMs, such as Stable Diffusion (v1.5, v2.1, v3.0) [47, 11], SDXL [45], both of which were pretrained on extensive text-to-image (T2I) datasets (i.e., LAION-5B [51]), have fueled progress across generative domains. These models offer powerful priors for image synthesis, and have been widely adapted through fine-tuning, particularly for natural image applications [68, 73, 26, 15]. Leveraging their strong generative capacity, we adapt a pretrained LDM for the SET task.
@@ 3.1 Overview of Proposed C-DiffSET
We utilize a paired SAR-EO dataset =$\mathcal{I}=\{(\mathbf{X},\mathbf{Y})\}$, where $\mathbf{X}\in\mathbb{R}^{C_{\text{sar}}\times H\times W}$ represents SAR images of $H\times W$ sizes and $C_{\text{sar}}$ channels and $\mathbf{Y}\in\mathbb{R}^{C_{\text{eo}}\times H\times W}$ denotes EO images of $H\times W$ sizes and $C_{\text{eo}}$ channels. Our goal is to generate a predicted EO image $\widehat{\mathbf{Y}}$ corresponding to the given SAR input $\mathbf{X}$. Fig. 3 illustrates the proposed Confidence Diffusion for SAR-to-EO Translation (C-DiffSET) framework, which comprises three key components: (i) the embedding of SAR and EO images into the latent space, (ii) the forward diffusion process, and (iii) the reverse diffusion process with the confidence-guided diffusion (C-Diff) loss. As described in Sec. 3.2, SAR image $\mathbf{X}$ and EO image $\mathbf{Y}$ are passed through the VAE encoder $\mathcal{E}_{\text{vae}}$, generating the latent features $\mathbf{z}_{x}\in\mathbb{R}^{C\times h\times w}$ (SAR feature) and $\mathbf{z}_{y}\in\mathbb{R}^{C\times h\times w}$ (EO feature). In the forward diffusion process (Sec. 3.3), noise $\bm{\epsilon}$ is added to the EO feature $\mathbf{z}_{y}$ over timesteps $t$s. During the reverse diffusion process, the U-Net $\psi$ takes the noisy EO feature $\mathbf{z}_{y}^{t}$, the timestep $t$, and the SAR feature $\mathbf{z}_{x}$ as conditional inputs. The Denoising U-Net $\psi$ predicts both the noise $\hat{\bm{\epsilon}}_{t}$ and a confidence map $\hat{\mathbf{c}}_{t}$ that quantifies pixel-wise uncertainty of the prediction. Finally, in Sec. 3.4, we describe the inference stage, where the predicted EO image $\widehat{\mathbf{Y}}$ is generated by reversing the noise-added process.
@@ 3.2 SAR and EO Latent Space Generation
EO latent space. Since EO images $\mathbf{Y}$ are naturally represented as 3-channel RGB inputs, they are directly passed through the VAE without modification. In our experiments, we confirmed that the reconstruction error $\|\mathbf{Y}-\mathcal{D}_{\text{vae}}\left(\mathcal{E}_{\text{vae}}\left(\mathbf{Y}\right)\right)\|_{2}$ is minimal, ensuring that the VAE accurately preserves the content and structure of EO images. This indicates that the VAE’s output provides an upper bound on the achievable performance of our framework, serving as a stable baseline for EO image generation.
SAR latent space. SAR data $\mathbf{X}$, depending on the satellite sensors, is available as single-polarization SAR images of 1-channel (HH or VV component) or full-polarization ones of 4-channel (HH, HV, VH, and VV components). For 1-channel SAR images, we repeat the channel three times to match the VAE’s input requirements. For 4-channel SAR data, we construct a 3-channel input by using HH, the average of HV and VH, and VV components. After passing these inputs through the VAE, we observed that the reconstruction error $\|\mathbf{X}-\mathcal{D}_{\text{vae}}\left(\mathcal{E}_{\text{vae}}\left(\mathbf{X}\right)\right)\|_{2}$ remained low, and the outputs appeared visually pleasing, as shown in Fig. 4. Additionally, we examined the VAE’s behavior under varying levels of speckle noise added to SAR inputs. The results demonstrate that the VAE’s reconstruction process adapts spatially to the noise level, effectively denoising the inputs while preserving structural details. This suggests that the VAE can embed SAR images into the same latent space as EO images without additional training, allowing SAR features to serve as conditioning inputs during the diffusion process. This ensures that pixel-wise correspondence between SAR and EO latent spaces is maintained throughout the generation process.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Satellite imagery plays a critical role in various applications, including surveillance, transportation, agriculture, disaster assessment, and environmental monitoring [84, 60, 71, 4, 68, 52, 42, 12]. A significant portion of these applications relies on Electro-Optical (EO) imagery, which provides multi-spectral data captured by EO sensors. However, a fundamental limitation of EO imagery is its susceptibility to weather and lighting conditions, reducing its usability in scenarios such as cloudy weather or nighttime. In contrast, Synthetic Aperture Radar (SAR) offers robust sensing capabilities under all weather conditions without the need for light, making it ideal for nighttime operations. Despite these advantages, SAR imagery is generally very difficult to directly interpret the structural and contextual information due to the very different nature of its image structures with high speckle noises [13, 8, 81, 57], unlike the natural color images. Therefore, for easy interpretation and downstream tasks such as target detection and recognition [40, 41, 39, 62, 14] developed for natural color images or satellite EO-RGB images, the SAR-to-EO image translation (SET) [76, 50, 25, 31, 32, 54, 15, 28] has been demanded for the usability expansion for SAR image applications. In SET, most existing approaches, including ours, focus on generating RGB bands of EO images [76, 50, 25, 54, 15, 28], as RGB representations are widely used for both human perception and vision-based deep learning models. Therefore, the SET problem in this paper is focused on generating the RGB bands of EO images for SAR input images.
@@
Recent advancements [76, 50, 25, 31, 32, 54, 15, 28] have demonstrated the potential of generating EO-like outputs from SAR images, offering improved textures and enhanced color information. However, several challenges persist: (i) The domain gap between SAR and EO images makes SET an ill-posed problem; (ii) The misalignment between SAR and EO datasets is prone to occur due to their different sensor platforms, satellite positioning shifts, or acquisition conditions, as shown in Fig. 2; (iii) Temporal discrepancies occur due to different acquisition times, also resulting in the spatial misalignment with differences in object appearance and seasons. This makes it very complicated the SET problem (Fig. 2); (iv) Finally, the scarcity of paired SAR-EO datasets makes it challenging for existing methods to learn the SET effectively, often leading to overfitting or unstable results.
@@
To address these limitations, we firstly propose a novel framework that leverages a pretrained Latent Diffusion Model (LDM) [47, 11] to improve SAR-to-EO image translation, which is denoted as Confidence Diffusion for SAR-to-EO Translation (C-DiffSET). The advantages of utilizing the pretrained LDM for SET tasks are as follows: (i) Since RGB bands of EO images share visual characteristics with natural images, we fine-tune the pretrained LDM—trained on large-scale natural image datasets—to transfer its representation power to the SET task. This addresses the issue of limited SAR-EO paired data, as our framework can leverage the pretrained knowledge from natural image distributions; (ii) The pretrained LDM operates in a 1/8-downsampled latent space, inherently alleviating spatially local misalignments caused by imperfect alignment processes; (iii) The variational auto-encoder (VAE) [29] from LDM can also be applicable for SAR images. As shown in Fig. 4, despite the presence of significant speckle noise in SAR data, the VAE encoder effectively embeds SAR images into the same latent space as EO images, leveraging its denoising capability. This enables the smooth transfer of SAR latents as conditioning information in the reverse diffusion process, ensuring the generation of EO outputs with accurate pixel-wise correspondence to the SAR inputs.
@@
Figure 2: Examples of misalignments and discrepancies in paired SAR-EO datasets. Left: Local spatial misalignments caused by sensor differences or acquisition conditions. Right: Temporal discrepancies where objects (e.g., ships) appear or disappear between SAR and EO images due to their different acquisition times.
@@
In addition to utilizing the pretrained LDM as a foundation model for our SET task, we introduce a confidence-guided diffusion (C-Diff) loss to handle temporal discrepancies. The temporal discrepancies, such as objects appearing in only one modality, can introduce artifacts and hallucinated content in the generated EO images. Since these discrepancies originate from real-world acquisition conditions that SAR and EO images are often obtained for same regions at different time instances, they cannot be explicitly corrected in the dataset itself during the training process. To mitigate this, the U-Net [48] in our framework predicts both each predicted noise and its corresponding confidence map, which quantifies the pixel-wise uncertainty in the predicted noise. The U-Net receives a channel-wise concatenated SAR and noisy EO features as input, allowing it to jointly model both contexts. This setup enables the confidence map to reflect the inherent uncertainty caused by temporal discrepancies, guiding the C-Diff loss to adaptively reduce penalties in regions where SAR-EO discrepancies likely occur. As a result, the generated EO images achieve high pixel-wise fidelity while minimizing artifacts and hallucinations. Our key contributions are summarized as:
@@
  • We propose C-DiffSET, the first framework to fine-tune a pretrained LDM for SET tasks, effectively leveraging their learned representations to overcome the scarcity of SAR-EO image pairs. In the C-DiffSET, SAR and EO images are embedded into the same latent space to ensure pixel-wise correspondence between SAR and EO latent throughout the framework;
@@
  • We introduce a novel C-Diff loss that can guide our C-DiffSET to reliably predict both EO outputs and confidence maps for accurate SET to mistigate locally mismatching challenges from temporal discrepancy due to their different acquisition times;
@@ 2.1 Image-to-Image Translation
Image-to-image translation has been widely studied in various fields, such as image colorization [78, 79], style transfer [87, 24, 55], and image inpainting [22, 44]. These models typically rely on carefully constructed paired datasets across domains for effective training. For instance, Pix2Pix [24] introduced conditional GANs with an $l_{1}$ loss for domain-specific translations. However, acquiring paired datasets remains challenging, and various methods have been proposed to work around unpaired data constraints [1, 35, 27, 7, 20, 6]. CycleGAN [87] addressed this limitation using a cycle-consistency loss to facilitate training without paired data. Further addressing non-bijective translation, StegoGAN [70] introduced steganography into GAN-based models, enhancing semantic consistency in cases of domain mismatch by reducing spurious features in generated images without additional supervision.
Diffusion-based approaches. Diffusion models [18, 47] have gained significant traction in image-to-image translation due to their ability to model complex data distributions through iterative denoising. Traditional Denoising Diffusion Probabilistic Models (DDPMs) [18] perform diffusion directly in the pixel space, but their high computational cost limits scalability in high-resolution tasks. To address this, Latent Diffusion Models (LDMs) [47] move the diffusion process to a learned latent space, significantly reducing memory and computational overhead while preserving high-resolution details. Palette [49] employs conditional DDPMs in the pixel domain for high-quality translations, whereas BBDM [33] performs diffusion directly in the latent space, eliminating the need for separate noise conditioning but imposing strict pixel-wise alignment requirements. DGDM [72] extends BBDM with a deterministic translator network to generate target features before diffusion, adding computational overhead and dependency on the translator’s performance. ControlNet [77] and its variants, Uni-ControlNet [82], extend pretrained diffusion models by injecting control signals through additional conditioning networks. These methods excel at preserving structural guidance for simple condition images, such as edge maps or skeleton poses. However, they are less suitable for the SET task, where the relationship between SAR and EO images is highly complex and requires learning intricate domain mappings rather than simple structural constraints.
@@ 2.2 SAR-to-EO Image Translation (SET)
Deep learning-based image-to-image translation methods [24, 55, 22, 20, 1, 35, 85, 33] have been widely adopted to remote sensing applications, including SAR-to-EO image translation (SET). However, SAR imagery poses unique challenges due to inherent characteristics like speckle noise from backscattering effects [13, 8, 81, 57, 59, 75, 21], complicating the translation process. To address these challenges, most SET studies employ GAN-based approaches [64, 5, 67, 32, 30, 9, 76, 73, 25, 34, 52]. Conditional GAN-based models are commonly used to leverage noisy SAR information as a conditioning input, while CycleGAN-based models align the SAR and EO domains by enforcing cycle-consistency constraints. SAR-SMTNet [73] introduced a Swin-Transformer-based [36] generator to improve structural consistency in SAR-to-EO translation. Additionally, CFCA-SET [31] proposed a coarse-to-fine SAR-to-EO model that incorporates near-infrared (NIR) images during training to refine EO generation.
@@
Despite these advancements, both GAN-based and diffusion-based methods encounter limitations stemming from the scarcity of paired SAR-EO datasets. The GAN-based models frequently face with convergence challenges and mode collapse under limited data, while diffusion-based approaches struggle with overfitting and slow convergence due to extensive training from scratch. Recent open-source [63] releases of foundation LDMs, such as Stable Diffusion [47, 11] and SDXL [45], pretrained on large-scale text-to-image datasets (e.g., LAION-5B [51]), have demonstrated strong generative priors for image synthesis. These models have been widely adapted for natural image applications [69, 74, 26, 16], showcasing their flexibility in domain adaptation. Building on this foundation, we fine-tune a pretrained LDM for the SET task. By using the extensive visual representations learned from large-scale generative models, C-DiffSET effectively adapts to SAR-EO data, achieving robust translation while mitigating the risk of overfitting.
@@ 3.1 Overview of Proposed C-DiffSET
We utilize a paired SAR-EO dataset =$\mathcal{I}=\{(\mathbf{X},\mathbf{Y})\}$, where $\mathbf{X}\in\mathbb{R}^{H\times W\times C_{\text{sar}}}$ represents SAR images of $H\times W$ sizes and $C_{\text{sar}}$ channels and $\mathbf{Y}\in\mathbb{R}^{H\times W\times C_{\text{eo}}}$ denotes EO images of $H\times W$ sizes and $C_{\text{eo}}$ channels. Our goal is to generate a predicted EO image $\widehat{\mathbf{Y}}$ corresponding to the given SAR input $\mathbf{X}$. Fig. 3 illustrates the proposed Confidence Diffusion for SAR-to-EO Translation (C-DiffSET) framework, which comprises three key components: (i) the embedding of SAR and EO images into the latent space, (ii) the forward diffusion process, and (iii) the reverse diffusion process with the confidence-guided diffusion (C-Diff) loss. As described in Sec. 3.2, SAR image $\mathbf{X}$ and EO image $\mathbf{Y}$ are passed through the VAE encoder $\mathcal{E}_{\text{vae}}$, generating the latent features $\mathbf{z}_{x}\in\mathbb{R}^{h\times w\times C}$ (SAR feature) and $\mathbf{z}_{y}\in\mathbb{R}^{h\times w\times C}$ (EO feature). In the forward diffusion process (Sec. 3.3), noise $\bm{\epsilon}$ is added to the EO feature $\mathbf{z}_{y}$ over timesteps $t$s. During the reverse diffusion process, the U-Net $\psi$ takes the noisy EO feature $\mathbf{z}_{y}^{t}$, the timestep $t$, and the SAR feature $\mathbf{z}_{x}$ as conditional inputs. The Denoising U-Net $\psi$ predicts both the noise $\hat{\bm{\epsilon}}_{t}$ and a confidence map $\hat{\mathbf{c}}_{t}$ that quantifies pixel-wise uncertainty of the prediction. Finally, in Sec. 3.4, we describe the inference stage, where the predicted EO image $\widehat{\mathbf{Y}}$ is generated by reversing the noise-added process.
@@ 3.2 SAR and EO Latent Space Generation
EO latent space. Since the RGB bands of EO images $\mathbf{Y}$ are naturally represented as 3-channel inputs, they are directly passed through the VAE without modification. In our experiments, we confirmed that the reconstruction error $\|\mathbf{Y}-\mathcal{D}_{\text{vae}}\left(\mathcal{E}_{\text{vae}}\left(\mathbf{Y}\right)\right)\|_{2}$ is minimal, ensuring that the VAE accurately preserves the content and structure of EO images. Furthermore, this low reconstruction error indicates that the RGB bands of EO images exhibit minimal domain gap compared to natural RGB images, making them well-suited for adaptation using pretrained LDMs. This indicates that the VAE’s output provides an upper bound on the achievable performance of our framework, serving as a stable baseline for EO image generation.
SAR latent space. SAR data $\mathbf{X}$, depending on the satellite sensors, is available as single-polarization SAR images of 1-channel (HH or VV component) or full-polarization ones of 4-channel (HH, HV, VH, and VV components). For 1-channel SAR images, we repeat the channel three times to match the VAE’s input requirements. For 4-channel SAR data, we construct a 3-channel input by using HH, the average of HV and VH, and VV components. After passing these inputs through the VAE, we observed that the reconstruction error $\|\mathbf{X}-\mathcal{D}_{\text{vae}}\left(\mathcal{E}_{\text{vae}}\left(\mathbf{X}\right)\right)\|_{2}$ remained low, and the outputs appeared visually pleasing, as shown in Fig. 4. Additionally, we examined the VAE’s behavior under varying levels of speckle noise added to SAR inputs. The results demonstrate that the VAE’s reconstruction process adapts spatially to the noise level, effectively denoising the inputs while preserving structural details. This indicates that the VAE can embed SAR images into the same latent space as EO images without additional training, allowing SAR features to serve as conditioning inputs during the diffusion process. This ensures that pixel-wise correspondence between SAR and EO latent spaces is maintained throughout the generation process.